OpenStudy (anonymous):

Subtract. –0.61 – (–7.74)

OpenStudy (anonymous):

@Thebossofme191

jabez177 (jabez177):

Use a calculator. :)

OpenStudy (thebossofme191):

7.13

OpenStudy (animallover456):

-7.13 is your answer because - - = -

OpenStudy (animallover456):

can i have medal

OpenStudy (thebossofme191):

actually itz not use a calculator to solve it

OpenStudy (thebossofme191):

the answer iz 7.13

OpenStudy (animallover456):

NO!!!!!! the calculator on google won't give you a negative number...and if u do it on paper it comes out to -7.13 NOT 7.13

OpenStudy (thebossofme191):

brb

OpenStudy (animallover456):

thanks!

OpenStudy (sushi121212):

–0.61 – (–7.74)= -0.61+7.74= 7.74 -0.61 ______ 7.13

OpenStudy (animallover456):

-7.13 you need to drop down the negative!!!!

OpenStudy (sushi121212):

What negative? That's a subtraction sign.

OpenStudy (animallover456):

negative sign

OpenStudy (sushi121212):

A number subtracted by a negative number will get you a positive number, so –0.61 – (–7.74) will really equal -0.61+7.74, which subtracts to be positive 7.13.

OpenStudy (thebossofme191):

thank u